The Magellanic Clouds Newsletter

An electronic exchange on Magellanic Clouds research

Edited by Bryan C. Dunne Co-Editors Eva K. Grebel and You-Hua Chu

                                 mcnews@astro.uiuc.edu
                 http://www.astro.uiuc.edu/projects/mcnews/MCNews.shtml
                         http://www.astro.uni-bonn.de/~mcnews/

No. 67 November 13, 2002

Contents

Letter From The Editor  1
Abstracts of 9 Refereed Papers  2
Abstract of 1 Non-Refereed Paper 7
Recent astro-ph Listings  8

Letter From The Editor

Dear Colleagues,

The Magellanic Clouds Working Group is pleased to bring you issue 67 of the Magellanic Clouds Newsletter. This month, we present 10 submitted abstracts and a collection of recent astro-ph listings of interest to Magellanic Clouds researchers. As always, MC News is available in a variety of formats, including L^{A}TEX, PostScript, Adobe Acrobat (PDF), and HTML. To download your preferred format, please direct your web browser to http://www.astro.uiuc.edu/projects/mcnews/MCNews.shtml.

Best Wishes, Bryan Dunne Editor, MC News

Abstracts of Refereed Papers

The Magellanic Stream, High-Velocity Clouds and the Sculptor Group

Mary E. Putman^{1}, Lister Staveley-Smith^{2}, Kenneth C. Freeman^{3}, Brad K. Gibson^{4} and David G. Barnes^{4}

1 Center for Astrophysics and Space Astronomy, University of Colorado, Boulder, CO 80309-0389, USA; Hubble Fellow 2 Australia Telescope National Facility, CSIRO, P.O. Box 76, Epping, NSW 1710 Australia 3 Research School of Astronomy & Astrophysics, Australian National University, Weston Creek P.O., Weston, ACT, Australia 2611 4 Centre for Astrophysics & Supercomputing, Swinburne University, Mail #31, P.O. Box 218, Hawthorn, VIC, Australia 3122

The Magellanic Stream is a 100^{∘} × 10^{∘} filament of gas which lies within the Galactic halo and contains ∼ 2 × 10^{8} M_{⊙} of neutral hydrogen. In this paper we present data from the HI Parkes All Sky Survey (HIPASS) in the first complete survey of the ire Magellanic Stream and its surroundings. We also present a summary of the reprocessing techniques used to recover large-scale structure in the Stream. The substantial improvement in spatial resolution and angular coverage compared to previous surveys reveals a variety of prominent features, including: bifurcation along the main Stream filament; dense, isolated clouds which follow the entire length of the Stream; head-tail structures; and a complex filamentary web at the head of the Stream where gas is being freshly stripped away from the Small Magellanic Cloud and the Bridge. Debris which appears to be of Magellanic origin extends out to 20^{∘} from the main Stream filaments. The Magellanic Stream has a velocity gradient of 700 km s^{−1} from the Clouds to the tail of the Stream, ∼ 390 km s^{−1} greater than that due to Galactic rotation alone, therefore implying a non-circular orbit. The dual filaments comprising the Stream are likely to be relics from gas stripped separately from the Magellanic Bridge and the SMC. This implies: (a) the Bridge is somewhat older than conventionally assumed; and (b) the Clouds have been bound together for at least one or two orbits. The transverse velocity gradient of the Stream also appears to support long-term binary motion of the Clouds. A significant number of the most elongated cataloged Stream clouds (containing ∼ 1% of the Stream mass) have position angles aligned along the Stream. This suggests the presence of shearing motions within the Stream, arising from tidal forces or interaction with the tenuous Galactic halo. As previously noted, clouds within one region of the Stream, along the sightline to the less distant half (southern half on the sky) of the Sculptor Group, show anomalous properties. There are more clouds along this sightline than an y other part of the Stream and their velocity distribution significantly deviates from the gr adient along the Stream. We argue that this deviation could be due to a combination of halo material, and not to distant Sculptor clouds based on a spatial and kinematic comparison between the Sculptor Group galaxies and the anomalous clouds, and the lack of cloud detection in the northern half of the group. This result has significant implications for the hypothesis that there might exist distant, massive HVCs within the Local Group. Cataloged clouds within the Magellanic Stream do not have a preferred scale size. Their mass spectrum f(M_{HI}) ∝ M_{HI}^{−2.0} and column density spectrum f(N_{HI}) ∝ N_{HI}^{−2.8} are steep compared with Lyα absorbers and galaxies, and similar to the anomalous clouds along the Sculptor Group sightline. Comments: to appear in ApJ, Jan. 2003 WWW: http://casa.colorado.edu/˜ mputman

A New Look at the Large-Scale HI Structure of the LMC

L. Staveley-Smith, S. Kim, M. R. Calabretta, R. F. Haynes, M. J. Kesteven

We present a Parkes multibeam HI survey of the Large Magellanic Cloud (LMC). This survey, which is sensitive to spatial structure in the range 200 pc to 10 kpc, complements the Australia Telescope Compact survey, which is sensitive to structure in the range 15 pc to 500 pc. With an rms column density sensitivity of 8 x 10^{16} cm^{−2} for narrow lines and 4 x 10^{17} cm^{−2} for typical linewidths of 40 km s^{−1}, emission is found to be extensive well beyond the main body of the LMC. Arm-like features extend from the LMC to join the Magellanic Bridge and the Leading Arm, a forward counterpart to the Magellanic Stream. These features, whilst not as dramatic as those in the SMC, appear to have a common origin in the Galactic tidal field, in agreement with recent 2MASS and DENIS results for the stellar population. The diffuse gas which surrounds the LMC, particularly at pa’s 90 to 330 deg, appears to be loosely associated with tidal features, but loosening by the ram pressure of tenuous Galactic halo gas against the outer parts of the LMC cannot be discounted. High-velocity clouds, which lie between the Galaxy and the LMC in velocity and which appear in the UV spectra of some LMC stars, are found to be associated with the LMC if their heliocentric velocity exceeds about +100 km s^{−1}. They are possibly the product of energetic outflows from the LMC disk. The HI mass of the LMC is found to be (4.8±0.2) × 10^{8} M_{⊙} (for an assumed distance of 50 kpc), substantially more than previous recent measurements. Comments: Accepted by MNRAS WWW: http://www.atnf.csiro.au/˜ lstavele/papers/lmc mb paper.ps.gz e-mail: lister.staveley-smith@atnf.csiro.au

Far-Ultraviolet & Hα Spectroscopy of SNR0057−7226 in the SMC HII Region N66

Charles W. Danforth(1), Ravi Sankrit(1), William P. Blair(1), J. Christopher Howk(1) & You-Hua Chu(2)

(1) Johns Hopkins University; (2) University of Illinois at Urbana-Champaign

N66/NGC346 is the largest and brightest HII region in the Small Magellanic Cloud and contains at least one known supernova remnant, SNR0057−7226. Optical emission from the remnant is overwhelmed by the bright photoionized emission from the nebula, but the remnant has been detected by way of far ultraviolet absorption lines. Here we present data from the Far Ultraviolet Spectroscopic Explorer (FUSE) satellite showing strong OVI and CIII emission from a position at the edge of SNR0057−7226. We also present high-resolution, long-slit Hα spectra across N66 showing high- and low-velocity emission corresponding closely to the X-ray boundaries of the supernova remnant. We use these FUV and optical data to determine the physical parameters of the shock and interaction geometry with N66. We find that ionizing photons from the many massive cluster stars nearby likely affect the ionization balance in the post-shock gas, hindering the production of lower-ionization and neutral species. We discuss the importance and inherent difficulty of searching for supernova remnants in or near bright HII regions and suggest that the far ultraviolet provides a viable means to discover and study such remnants. Comments: Submitted to the Astrophysical Journal WWW: http://fuse.pha.jhu.edu/˜ danforth/n66/snrpaper.ps e-mail: danforth@pha.jhu.edu

X-Ray and Optical Observations of XTE J0052-723 a Transient Be/X-Ray Pulsar in the Small Magellanic Cloud

S. Laycock (1), R. H. D. Corbet (2,3), M. J. Coe (1), F. E. Marshall (2), C. Markwardt (2,4), W. Edge (1)

(1) Department of Physics and Astronomy, Southampton University SO17 1BJ

(2) NASA Goddard Space Flight Center, Greenbelt, MD 20771 USA

(3) Universities Space Research Association

(4) Department of Astronomy, University of Maryland, College Park, MD 20742, USA

On December 27th 2000 during our regular SMC monitoring program with Rossi X-ray Timing Explorer strong pulsations were detected with a period of 4.78 seconds. Subsequent slew observations performed on Jan 9th and 13th across the field of view allowed localisation of the pulsar’s position to RA: 00 52 17, Dec: 72 19 51 (J2000). The outburst continued until Jan 24th, 7 PCA observations were obtained during this period, yielding a maximum X-ray luminosity ∼10^{38} erg s^{−1}. Following calculation of the pulsar position, optical observations of the RXTE error box were made on Jan 16th 2001 with the 1m telescope of the South African Astronomical Observatory (SAAO) while the source was still in X-ray outburst. Candidate Be stars identified from their photometric colours were subsequently observed with the SAAO 1.9m on Nov 7th 2001 to obtain spectra. Only one of the photometrically identified stars [MA93]537 showed prominent Hα emission, with a double peaked line-profile (EW= -43.3+/-0.7 Å, separation velocity=200+/-15 km s^{−1}) confirming the presence of a substantial circumstellar disk. Comments: Accepted for publication in MNRAS WWW: http://arXiv.org/abs/astro-ph/0210397 e-mail: sgtl@astro.soton.ac.uk

AGB stars in the Magellanic Clouds. I. The C/M ratio.

M.-R.L. Cioni (1) & H.J. Habing (2)

(1) ESO, Karl-Schwarzschild-str. 2, D-85748 Garching bei Muenchen, Germany

(2) Sterrewacht Leiden, Niels Bohrweg 2, 2333 RA Leiden, The Netherlands

Regions of different metallicity have been identified in the Magellanic Clouds by using the ratio between Asymptotic Giant Branch stars of spectral type C and M. In the Large Magellanic Cloud the ratio appears to decrease radially while in the Small Magellanic Cloud (SMC) there is no clear trend, reflecting either the large extension of the SMC along the line of sight or a more complex star formation history. The distribution of the C/M ratio is clumpy and corresponds to a spread in [Fe/H] of 0.75 dex in both Clouds. There is an indication of increasing C/M ratio, thus decreasing metallicity, towards the Bridge region connecting the two Clouds. Comments: 8 pages, 8 figures, submitted to A&A e-mail: mcioni@eso.org

Testing intermediate-age stellar evolution models with VLT photometry of LMC clusters. I. The data.

Carme Gallart(2,3), Manuela Zoccali(5), Giampaolo Bertelli(6,9), Cesare Chiosi(7), Pierre Demarque(4), Leo Girardi(8), Emma Nasi(9), Jong-Hak Woo(4), Sukyoung

Yi(10)

(1) Based on observations collected at the European Southern Observatory, Paranal, Chile (ESO N^{o} 64.L-0385)

(2) Andes Prize Fellow, Universidad de Chile and Yale University

(3) Currently: Ramón y Cajal Fellow. Instituto de Astrofísica de Canarias, 38200 Tenerife, Canary Islands, Spain; carme@iac.es

(4) Department of Astronomy, Yale University, P.O. Box 208101, New Haven, CT 06520-8101; demarque@astro.yale.edu, jhwoo@astro.yale.edu

(5) European Southern Observatory, Karl Schwarzschild Strasse 2, D-85748 Garching bei München, Germany; mzoccali@eso.org

(6) National Council of Research, IAS-CNR, Rome, Italy; bertelli@pd.astro.it

(7) Dipartimento di Astronomia dell’Università di Padova, Vicolo dell’Osservatorio 5, I-35122 Padova, Italy; chiosi@pd.astro.it (8) Osservatorio Astronomico di Trieste, Via Tiepolo 11, I-34131 Trieste, Italy girardi@pd.astro.it

(9) Osservatorio Astronomico di Padova, Vicolo dell’Osservatorio 5, 35122 Padova, Italy; nasi@pd.astro.it

(10) Astrophysics, University of Oxford, Keble Road, Oxford, OX1 3RH, United Kingdom; yi@astro.ox.ac.uk

This is the first of a series of three papers devoted to the calibration of a few parameters of crucial importance in the modeling of the evolution of intermediate-mass stars, with special attention to the amount of convective core overshoot. To this end we acquired deep V and R photometry for three globular clusters of the Large Magellanic Cloud (LMC), namely NGC 2173, SL 556 and NGC 2155, in the age interval 1–3 Gyr. In this first paper, we describe the aim of the project, the VLT observations and data reduction, and we make preliminary comparisons of the color-magnitude diagrams with both Padova and Yonsei-Yale isochrones. Two following papers in this series present the results of a detailed analysis of these data, independently carried out by members of the Yale and Padova stellar evolution groups. This allows us to compare both sets of models and discuss their main differences, as well as the systematic effects that they would have to the determination of the ages and metallicities of intermediate-age single stellar populations. Comments: AJ, in press WWW: http://xxx.lanl.gov/abs/astro-ph/0208188 e-mail: carme@iac.es

Testing intermediate-age stellar evolution models with VLT photometry of LMC clusters. II. Analysis with the Yale models.

J.-H. Woo (Yale), C. Gallart (Univ. de Chile), P. Demarque (Yale), S. Yi (Oxford), M. Zoccali (ESO)

We present an analysis of CMDs of three intermediate-age LMC clusters, namely NGC 2173, SL 556 and NGC 2155. The main goal of our project is to investigate the amount of convective core overshoot necessary to reproduce the CMDs of relatively metal-poor, intermediate age stellar populations. We conclude that a moderate amount of overshoot and some fraction of binary stars are essential for reproducing the observed shapes around the turnoff in the CMD’s of all three clusters: unresolved binary stars fill in the expected core contraction gap, and make a unique sequence near the gap, which cannot be reproduced by single stars alone, even with a larger amount of overshoot. From our overall analysis such as, shape of isochrones, star counts, color distribution, and synthetic CMD comparisons, we conclude that overshoot ∼20% of the local pressure scale height best reproduces the CMD properties of all three clusters. The best age estimates are 1.5, 2.1 and 2.9 Gyr for NGC 2173, SL 556 and NGC 2155, respectively. Comments: AJ, in press WWW: http://xxx.lanl.gov/abs/astro-ph/0208142 e-mail: carme@iac.es

Testing intermediate-age stellar evolution models with VLT photometry of LMC clusters. III. Padova results

G. Bertelli (CNR, OAPd), E. Nasi (OAPd), L. Girardi (OATs), C. Chiosi (ADPd), M. Zoccali (ESO), C. Gallart (IAC)

The color-magnitude diagrams (CMDs) of three intermediate-age LMC clusters, NGC 2173, SL556 and NGC2155 are analyzed to determine their age and metallicity basing on Padova stellar models. Synthetic CMDs are compared with cluster data. The best match is obtained using two fitting functions based on star counts in the different bins of the cluster CMD. Two different criteria are used. One of them takes into account the uncertainties in the color of the red clump stars. Given the uncertainties on the experimental values of the clusters metallicity, we provide a set of acceptable solutions. They define the correspondent values of metallicity, age, reddening and distance modulus (for the assumed IMF). The comparison with Padova models suggests for NGC 2173 a prolonged star formation (spanning a period of about 0.3 Gyr), beginning 1.7 Gyr and ending 1.4 Gyr ago. The metallicity Z is in the range 0.0016 − 0.003. Contrary to what suggested for NGC 2173 a period of extended star formation was not required to fit the SL 556 and the NGC 2155 observations. For SL 556 an age of 2.0 Gyr is obtained. The metallicity value is in the range 0.002 − 0.004, depending on the adopted comparison criterium. The derived age for NGC 2155 is 2.8 Gyr and its metallicity Z is in the range 0.002 − 0.003. The CMD features of this cluster suggest that a more efficient overshoot should be adopted in the evolutionary models. Comments: Accepted by the Astronomical Journal WWW: http://arXiv.org/abs/astro-ph/0211169 e-mail: nasi@archimede.pd.astro.it

The Metallicity-Luminosity Relation, Effective Yields, and Metal Loss in Spiral and Irregular Galaxies

Donald R. Garnett

Steward Observatory, University of Arizona

I present results on the correlation between galaxy mass, luminosity, and metallicity for a sample of spiral and irregular galaxies having well-measured abundance profiles, distances, and rotation speeds. Additional data for low surface brightness galaxies from the literature are also included for comparison. These data are combined to study the metallicity-luminosity and metallicity-rotation speed correlations for spiral and irregular galaxies. The metallicity luminosity correlation shows its familiar form for these galaxies, a roughly uniform change in the average present-day O/H abundance of about a factor 100 over 11 magnitudes in B luminosity. However, the O/H - V_{rot} relation shows a change in slope at a rotation speed of about 125 km s^{−1}. At faster V_{rot}, there appears to be no relation between average metallicity and rotation speed. At lower V_{rot}, the metallicity correlates with rotation speed. This change in behavior could be the result of increasing loss of metals from the smaller galaxies in supernova-driven winds. This idea is tested by looking at the variation in effective yield, derived from observed abundances and gas fractions assuming closed box chemical evolution. The effective yields derived for spiral and irregular galaxies increase by a factor of 10-20 from V_{rot} ≈ 5 km s^{−1} to V_{rot} ≈ 300 km s^{−1}, asympotically increasing to approximately constant y_{eff} for V_{rot} greater thanabout 150 km s^{−1}. The trend suggests that galaxies with V_{rot} less than about 100-150 km s^{−1} may lose a large fraction of their SN ejecta, while galaxies above this value tend to retain metals. Comments: Accepted for publication in the Astrophysical Journal (20 December 2002 issue) WWW: http://ocotillo.as.arizona.edu/˜ dgarnett/Papers/metlum.ps e-mail: dgarnett@as.arizona.edu

Abstracts of Non-Refereed Papers

Element Abundances in Nearby Galaxies

Donald R. Garnett

Steward Observatory, University of Arizona

In these lectures I present a highly opinionated review of the observed patterns of metallicity and element abundance ratios in nearby spiral, irregular, and dwarf elliptical galaxies, with connection to a number of astrophysical issues associated with chemical evolution. I also discuss some of the observational and theoretical issues associated with measuring abundances in H II regions and gas and stellar surface densities in disk galaxies. Finally, I will outline a few open questions that deserve attention in future investigations.

Comments: Invited review for the XIIIth Canary Islands Winter School on Astrophysics, ”Cosmochemistry: the Melting Pot of the Elements” WWW: http://ocotillo.as.arizona.edu/˜ dgarnett/Papers/garnettws.ps e-mail: dgarnett@as.arizona.edu

Recent astro-ph Listings

astro-ph/0211169:
Title: Testing intermediate-age stellar evolution models with VLT photometry of LMC
clusters. III. Padova results
Authors: G. Bertelli (CNR, OAPd), E. Nasi (OAPd), L. Girardi (OATs), C. Chiosi
(ADPd), M. Zoccali (ESO), C. Gallart (IAC)
Comments: 41 pages with 7 figures. Accepted by the Astronomical Journal. Paper I
and II appeared as astro-ph/0208188 and astro-ph/0208142
astro-ph/0211153:
Title: Chemical evolution and nature of Damped Lyman-Alpha systems
Authors: F. Calura (1), F. Matteucci (1), G. Vladilo (2) ((1) Dip. Astronomia, Universita'
di Trieste; (2) INAF, Osservatorio Astronomico, Trieste)
Comments: MNRAS, accepted for publication, 19 pages, 14 figs
astro-ph/0211060:
Title: The disruption time of clusters in selected regions of four galaxies
Authors: Henny J.G.L.M. Lamers, Stratos G. Boutloukos
Comments: 6 pages, 8 figures. To be published in ESO Symposia: "Extragalactic
Globular Cluster Systems", ed. by M. Kissler-Patig
astro-ph/0211047:
Title: Dynamical Evolution of Globular Clusters in the Magellanic Clouds and Other
Local Systems
Authors: A.D. Mackey, G.F. Gilmore (IoA, Cambridge, UK)
Comments: 2 pages, 2 figures, to be published in ``New Horizons in Globular Cluster
Astronomy'', ASP Conference Series, 2002, eds. G. Piotto, G. Meylan, G. Djorgovski,
M. Riello
astro-ph/0211029:
Title: High-velocity regions in planetary nebulae
Authors: Krzysztof Gesicki, Albert A. Zijlstra
Comments: accepted for publication in MNRAS
astro-ph/0210650:
Title: Multiwavelength Time Series Data of the LMC Cluster Reticulum
Authors: M. Monelli (1), G. Andreuzzi (2), G. Bono (2), R. Buonanno (1), F. Caputo (2),
M. Castellani (2), V. Castellani (2), C. E. Corsi (2), M. Dall' Ora (1), G. Marconi (3), L.
Pulone (2), V. Ripepi (4), J. Storm (5), V. Testa (2) ((1) Universita' di Roma Tor
Vergata, (2) INAF-Osservatorio Astronomico di Roma, (3) ESO-Chile, (4)
INAF-Osservatorio Astronomico di Napoli, (5) Postdam Observatory)
Comments: 2 pages, 1 figure, uses newpasp.sty, to be published in ``New Horizons in
Globular Cluster Astronomy'', ASP Conference Series, 2002; Eds.: G. Piotto, G.
Meylan, G. Djorgowski and M. Riello
astro-ph/0210615:
Title: High Resolution \hi observations of the Western Magellanic Bridge
Authors: E. Muller, L. Staveley-Smith., W. Zealey, S. Stanimirovic
Comments: 23 pages, 15 figures. Accepted for publication by MNRAS. For paper with
higher resolution figures: this http URL
astro-ph/0210595:
Title: Star Cluster Formation and Disruption Time-Scales -- I. An empirical
determination of the disruption time of star clusters in four galaxies
Authors: S.G. Boutloukos, H.J.G.L.M. Lamers
Comments: 18 pages, 18 figures. Accepted for publication by Monthly Notices

astro-ph/0210563:
Title: Study of a Strategy for Parallax Microlensing Detection Towards the Magellanic
Clouds
Authors: S. Rahvar (Sharif University of Technology, Tehran, Iran), M. Moniez
(IN2P3-CNRS, France), R. Ansari (IN2P3-CNRS, France), O. Perdereau
(IN2P3-CNRS, France)
Comments: Comments: 10 pages, 12 figures. Submitted to A&A
astro-ph/0210549:
Title: Kinematic outliers in the LMC: constraints on star-star microlensing
Authors: HongSheng Zhao (1), Rodrigo Ibata (2), Geraint Lewis (3,4), Mike Irwin (1)
((1) IoA Cambridge, (2) Obs. Strabourg, (3) AAO, (4) University of Sydney)
Comments: 6pp w/ 10 eps fig (MNRAS accepted)
astro-ph/0210501:
Title: A New Look at the Large-Scale HI Structure of the LMC
Authors: L. Staveley-Smith, S. Kim, M. R. Calabretta, R. F. Haynes, M. J. Kesteven
Comments: Accepted by MNRAS, 19 pages, 13 figurs, 21 eps files, full resolution
paper (1.4 M) available at this http URL
astro-ph/0210498:
Title: Pulsation at the tip of the first giant branch?
Authors: Yoshifusa Ita (1) Toshihiko Tanab (1), Noriyuki Matsunaga (1), Yasushi
Nakajima (2), Chie Nagashima (2), Takahiro Nagayama (2), Daisuke Kato (2), Mikio
Kurita (2), Tetsuya Nagata (2), Shuji Sato (2), Motohide Tamura (3), Hidehiko Nakaya
(4), Yoshikazu Nakada (1,5) ((1) Institute of Astronomy, University of Tokyo; (2)
Department of Astrophysics, Nagoya University; (3) National Astronomical
Observatory of Japan; (4) Subaru Telescope, National Astronomical Observatory of
Japan; (5) Kiso Observatory, School of Science, The University of Tokyo)
Comments: 5 pages, 3 figures; accepted for publication in MNRAS Letters
astro-ph/0210397:
Title: X-Ray and Optical Observations of XTE J0052-723 a Transient Be/X-Ray
Pulsar in the Small Magellanic Cloud
Authors: S. Laycock (1) R. H. D. Corbet (2,3), M. J. Coe (1), F. E. Marshall (2), C.
Markwardt (2,4), W. Edge (1). ((1)Southampton University,(2)NASA/GSFC,
(3)USRA,(4)University of Maryland)
Comments: 8 pages, 8 figures, Accepted for publication in MNRAS
astro-ph/0210380:
Title: The intriguing case of Sk-69 194
Authors: Pablo G. Ostrov
Comments: 5 pages, 3 postscript figures. Submitted to the MNRAS
astro-ph/0210367:
Title: The Optical Gravitational Lensing Experiment. Is Interstellar Extinction Toward
the Galactic Center Anomalous?
Authors: A. Udalski
Comments: 10 pages
astro-ph/0210295:
Title: Ten eclipsing binaries in the Small Magellanic Cloud: fundamental parameters
and SMC distance
Authors: Tim J. Harries, Ron W. Hilditch, Ian D. Howarth
Comments: 18 pages, accepted by MNRAS
astro-ph/0210241:
Title: Long-term Properties of Accretion Disks in X-ray Binaries: I. the variable third
period in SMC X-1

Authors: W.I. Clarkson, P. A. Charles, M. J. Coe, S.G.T. Laycock, M.D. Tout, C.A.
Wilson
Comments: 6 figures Re-submitted to MNRAS after referee's comments
astro-ph/0210225:
Title: On the origin of nitrogen
Authors: L.S. Pilyugin (Main Astron. Obs. of Nat. Academy of Sc. of Ukraine), T.X.
Thuan (University of Virginia, Charlottesville, VA), J.M. Vilchez (Instituto de
Astrofisica de Andalucia, Granada, Spain)
Comments: 16 pages, 11 figures, accepted for publication in Astronomy and
Astrophysics
astro-ph/0210157:
Title: A Grid of NLTE Line-Blanketed Model Atmospheres of O-type Stars
Authors: Thierry Lanz, Ivan Hubeny
Comments: 54 pages, inluding 14 figures; submitted to Astrophysical Journal
astro-ph/0210060:
Title: Cluster formation in the LMC
Authors: A. Vallenari, A. Moretti, E.V. Held, L.Rizzi, D.Bettoni
Comments: 2 pages, Proceedings of the Conference : New Horizons in Globular
Clusters